# A multiple imputation approach to distinguish curative from life-prolonging effects in the presence of missing covariates

### v1

@@ 4.1 Simulation design
We conducted a simulation study to compare the performance of the imputation methods for cure models presented in Section 3. The comparison was carried out considering several scenarios, defined as a function of the missing data percentages, model parameters, and variables’ distributions.
In this simulation study, we generated 1000 datasets, each containing 500 observations. The variables used in the study are $W$, $X$ and $Z$. Both $X$ and $Z$ follow a Bernoulli distribution with a probability of 0.5, while $W$ is either $Ber(0.5)$ or $N(0.5,1)$. The choice of a Normal variable with a mean of 0.5 rather than a standard Normal was made to ensure the stability of the cure rate when varying W’s distribution.
Variable $X$ is a covariate in the incidence model, $Z$ is a covariate in the latency model, and both of them have no missing values. Variable $W$ is a covariate in the models for both the incidence and the latency, and it is either missing completely at random (MCAR) with 15% or 30% missing values or missing at random (MAR) with 30% missing observations. To produce missing cases following a MAR mechanism, we used a multivariate amputation procedure (Schouten et al., 2018) in which the probability of $W$ to be missing depends on both $X$ and $Z$.
@@
$S(t)=\exp{(-\lambda t^{\rho}\exp{(\beta_{1}W+\beta_{2}Z)})}$
with $\lambda=0.25$, $\rho=1.45$, and truncated at 8, to satisfy the sufficient follow-up assumption. For cured individuals (having $G=0$), $T$ is set to an extremely large value, as they never experience the event of interest.
@@
The complete uncured indicator $G$ is generated as a Bernoulli random variable with probability $\pi$ defined by a logistic regression of the form
$P(G=1)=\frac{\exp{(\alpha_{0}+\alpha_{1}W+\alpha_{2}X)}}{1+\exp{(\alpha_{0}+\alpha_{1}W+\alpha_{2}X)}}$ (12)
In practice, we only use the complete indicator drawn from (12) for generating the follow-up times $Y$, namely to assign large values of $Y$ to cured individuals. Instead, for multiple imputation in the cure model, we use the observed uncured indicator, which is 1 for uncensored observations, 0 for observations in the plateau, and unknown otherwise.
We consider two sets of parameter values for the incidence and latency models. The first one is $\alpha_{0}=1$, $\alpha_{1}=-1$, $\alpha_{2}=0.5$, $\beta_{1}=-0.2$ and $\beta_{2}=0$. This set of parameters is based on the motivating application and therefore is only used when $W$ is Bernoulli distributed with around 15% of observations missing completely at random. This scenario leads to an average cure rate of $\sim 33$%, an average censoring rate of $\sim 45$%, and an average $\sim 18$% observations in the plateau.
The second set of parameters is $\alpha_{0}=0.1$ and $\alpha_{1}=\alpha_{2}=\beta_{1}=\beta_{2}=0.5$. These values are chosen to illustrate a more general setting. For this choice of parameters, we consider $W$ as Normally distributed with around 30% of observations missing at random or missing completely at random. Both the scenarios in which the missing mechanism is MCAR and MAR lead to an average cure rate of $\sim 36$%, an average censoring rate of $\sim 46$%, and an average $\sim 17$% observations in the plateau.
@@
  • Scenario A: MCAR data with 15% missingness; $W\sim$ Bernoulli; $\alpha_{0}=1$, $\alpha_{1}=-1$, $\alpha_{2}=0.5$, $\beta_{1}=-0.2$ and $\beta_{2}=0$.
  • Scenario B: MCAR data with 30% missingness; $W\sim$ Normal; $\alpha_{0}=0.1$ and $\alpha_{1}=\alpha_{2}=\beta_{1}=\beta_{2}=0.5$.
  • Scenario C: MAR data with 30% missingness; $W\sim$ Normal; $\alpha_{0}=0.1$ and $\alpha_{1}=\alpha_{2}=\beta_{1}=\beta_{2}=0.5$.
@@ 4.2 Simulation results
In Table 1, we present the results of simulations for the three different scenarios.
@@
Method $\alpha_{0}$ $\alpha_{1}$ $\alpha_{2}$ $\beta_{1}$ $\beta_{2}$
@@ 3 Empirical coverage of 95% confidence intervals
The performance of the imputation methods was assessed using different metrics, specifically the mean squared error (MSE), the width of 95% confidence intervals (CI width), and the empirical coverage of 95% confidence intervals (Coverage). The full data scenario provides a baseline for comparison, as it describes these metrics for models fitted to the original data, without missing values.
@@
In the scenario with 15% MCAR (Scenario A), the complete-case analysis, obtained by discarding units with missing values for $W$ and then fitting a cure model on the complete data only, exhibits slightly higher MSE and wider confidence intervals compared to the full data scenario. As expected, estimates from the complete case analysis are not biased but have a larger variance since fewer observations are being used. Both the exact and the approximate imputation methods performed similarly to the full data scenario, with only minor differences in CI widths, only for the estimates of the effects associated with $W$. In particular, we observe that multiple imputation lead to decreased variance for the estimates of the parameters corresponding to fully observed variables.
@@
When the missingness is increased to 30%, still under an MCAR mechanism (Scenario B), this trend is confirmed: the complete-case analysis shows higher MSE and wider confidence intervals, while the imputation methods (via either the exact or approximate procedure) yield results close to the full data scenario. Although there are slight variations, these methods effectively preserve empirical coverage, with a slightly greater CI length when the effects of $W$ are considered.
@@
These results show that the proposed approaches provide reliable parameter estimates even in the presence of a moderate to high proportion of missing data. However, the true strength of our methods becomes evident in the MAR scenario (Scenario C). Here, the proposed imputation methods significantly outperform the complete-case analysis according to all evaluation metrics. This demonstrates that when data are missing at random, both exact and approximate imputation methods can effectively mitigate the impact of missing data, producing estimates close to those obtained with full data. In contrast, the complete-case analysis performs poorly, exhibiting a marked increase in MSE and a decrease in coverage as a result of the bias. Specifically, for $\alpha_{0}$, the MSE of the complete-case approach increased to 0.35 and the coverage decreased to 0.32. Additionally, the CI width for $\alpha_{2}$ was considerably wider than that obtained with the imputation methods. Moreover, the estimate $\beta_{2}$ shows a slight deterioration compared to scenarios A and B. Indeed, when using complete-case analysis, the missing data of $W$ leads to the removal of observations that might be critical for accurately estimating $\alpha_{2}$ and $\beta_{2}$, especially if $W$ is correlated with $X$ and $Z$. This results in biased estimates and an increase in the MSE. In other words, by eliminating observations with missing $W$ data, the model’s ability to capture the true relationships involving $X$ and $Z$ may be reduced.
The exact conditional distribution method outperforms the approximate method in terms of CI width, indicating more precise estimates.
@@
Across all scenarios, both the exact and approximate methods consistently provide reliable estimates, maintaining the empirical coverage of the 95% confidence intervals. Particularly in cases where covariate $W$ is missing at random, the proposed methods significantly outperform the complete-case approach, demonstrating their ability to effectively handle complex missing data patterns.
@@ 4.3 Algorithm performance under model misspecification
Appendix A and B provide the exact and approximate distributions for cases where the variable $W$ is included in only the incidence or the latency of the Cox PH cure model. However, we may lack information about the impact of $W$ on the two components. Therefore, we aim to study the performance of the proposed imputation algorithm when the model is misspecified, e.g. when $W$ is present in only one of the two components and we assume it is in both. To do this, we conduct a simulation study, examining two scenarios:
  • Scenario D: MAR data with 30% missingness; $W\sim$ Normal; $\alpha_{0}=0.1$, $\alpha_{1}=\alpha_{2}=\beta_{2}=0.5$, and $\beta_{1}=0$
  • Scenario E: MAR data with 30% missingness; $W\sim$ Normal; $\alpha_{0}=0.1$, $\alpha_{2}=\beta_{1}=\beta_{2}=0.5$, and $\alpha_{1}=0$
These scenarios are similar to Scenario C introduced in section 4.1, but they differ in that the coefficient associated with the missing covariate $W$ is set to zero in either the latency component (Scenario D, $\beta_{1}=0$) or the incidence component (Scenario E, $\alpha_{1}=0$). This allows us to test the algorithm’s ability to handle situations where the missing variable affects only one component of the Cox PH cure model.
Importantly, our focus is testing the algorithm’s performance under the MAR scenario. We expect that if the algorithm performs well in these more complex conditions, it will likely do so in scenarios where the data is MCAR or the percentage of missing data is lower.
This simulation study provides insights into the robustness of the mixture cure model across the two distinct scenarios. If the imputation algorithm performs well under these misspecifications, we can derive that our model is robust and does not require prior knowledge of the covariate’s precise impact on incidence or latency. Instead, the model, through its fit to the available data, can suggest the appropriate placement of different covariates based on the significance of the coefficient estimates.
In Table 2, simulation results for scenarios E and D are shown. Analogously to what is seen in Table 1, the performance of the imputation methods was assessed using the MSE, the CI width, and the empirical coverage of 95% confidence intervals.
@@
Method $\alpha_{0}$ $\alpha_{1}$ $\alpha_{2}$ $\beta_{1}$ $\beta_{2}$
@@ 3 Empirical coverage of 95% confidence intervals
Quite similar to results in Table 1, the complete-case analysis shows a significant increase in MSE for $\alpha_{0}$, reaching 0.33 in Scenario D, with a corresponding coverage of only 0.31. This result reflects the method’s vulnerability to bias when missing data is not appropriately addressed. The complete-case method also displays increased CI widths, especially for $\alpha_{2}$, highlighting the loss of information and precision due to missing data.
The exact method generally maintains low MSE values in both scenarios and good coverage for most parameters. However, its performance sometimes falls short of the approximate method, as seen with $\alpha_{1}$ in Scenario D, where the exact method yields a coverage of 0.91 compared to 0.98 of the approximate method. The improved coverage of the approximate method can be attributed to its increased CI width, which provides a more conservative estimate, thereby encompassing more potential values of the parameter. The exact method’s performance suggests it may be more sensitive to model assumptions, leading to increased bias when these assumptions are not met. Despite this, the results support conclusions about the general robustness of the model. The slightly poorer performance compared to the approximate method could be due to the tighter linkage between the missing data variable $W$ and the linear predictors, resulting in higher sensitivity to model assumptions and less flexibility in handling deviations.
The approximate method shows consistent performance with low MSEs in both scenarios and high coverage, highlighting its adaptability in dealing with model deviations. It achieves slightly better coverage for parameters like $\beta_{1}$ in Scenario E, reaching 0.94 compared to the exact method’s 0.90.
In Scenario E, the complete-case analysis exhibits substantial MSE values for $\alpha_{0}$ and poor coverage, confirming its inadequacy in scenarios with high levels of missingness. Conversely, the approximate method maintains reliable coverage and CI widths across parameters due to its adaptability and reduced reliance on strict assumptions.
Overall, these results suggest that the approximate method is particularly well-suited for real-world applications where knowledge of model structure is limited and data are affected by several sources of error. Its robust performance across scenarios D and E reinforces the idea that it is a viable alternative to the exact method.

### v2

@@ 4.1 Simulation design
We conducted a simulation study to compare the performance of the imputation methods for cure models presented in Section 3. The comparison was carried out considering several scenarios, defined by varying the missing data percentages, model parameters, and variables’ distributions.
In this simulation study, we generated 1000 datasets, each containing 500 observations. The variables used in the study are $W$, $X$ and $Z$. Both $X$ and $Z$ follow a Bernoulli distribution with a probability of 0.5, while $W$ is either $Ber(0.5)$ or $N(0.5,1)$. The choice of a Normal variable with a mean of 0.5 rather than a standard Normal was made to ensure the stability of the cure rate when varying W’s distribution. The three variables have a correlation coefficient of 0.5.
Variable $X$ is a covariate in the incidence model, $Z$ is a covariate in the latency model, and both have no missing values. Variable $W$ is a covariate in the models for both incidence and latency, and it is either missing completely at random (MCAR) with 15% or 30% missing values or missing at random (MAR) with 30% missing observations. To produce a MAR mechanism, we used a multivariate amputation procedure (Schouten et al., 2018) in which the probability of $W$ to be missing depends on both $X$ and $Z$ and the outcome variable ($Y,\Delta)$.
@@
$S(t)=\exp{(-\lambda t^{\rho}\exp{(\beta_{W}W+\beta_{Z}Z)})}$
with $\lambda=0.25$, $\rho=1.45$, truncated at $t=8$, to satisfy the sufficient follow-up assumption. For cured individuals (having $G=0$), $T$ is set to an extremely large value, as they never experience the event of interest.
@@
The complete uncured indicator $G$ is generated as a Bernoulli random variable with probability $\pi$ defined by a logistic regression
$P(G=1)=\frac{\exp{(\alpha_{0}+\alpha_{W}W+\alpha_{X}X)}}{1+\exp{(\alpha_{0}+\alpha_{W}W+\alpha_{X}X)}}$ (12)
In practice, we only use the complete indicator drawn from (12) for generating the follow-up times $Y$, namely to assign large values of $Y$ to cured individuals. Instead, when performing multiple imputation in the cure model, we use the observed uncured indicator, which is 1 for uncensored observations, 0 for observations in the plateau, and unknown otherwise.
We consider two sets of parameter values for the incidence and latency models. The first one is $\alpha_{0}=1$, $\alpha_{W}=-1$, $\alpha_{X}=0.5$, $\beta_{W}=-0.2$ and $\beta_{Z}=0$. This set of parameters is based on the motivating application and therefore is only used when $W$ is Bernoulli distributed with around 15% of observations missing completely at random. This scenario leads to an average cure rate of $\sim 33$%, an average censoring rate of $\sim 45$%, and an average fraction of observations in the plateau of $\sim 18$%.
The second set of parameters is $\alpha_{0}=0.1$ and $\alpha_{W}=\alpha_{X}=\beta_{W}=\beta_{Z}=0.5$. These values are chosen to illustrate a more general setting. For this choice of parameters, we consider $W$ as Normally distributed with around 30% of observations missing at random or missing completely at random. Both the scenarios in which the missing mechanism is MCAR and MAR lead to an average cure rate of $\sim 36$%, an average censoring rate of $\sim 46$%, and an average fraction of observations in the plateau of $\sim 17$%.
@@
  • Scenario A: MCAR data with 15% missingness;
    $W\sim$ Bernoulli; $\alpha_{0}=1$, $\alpha_{W}=-1$, $\alpha_{X}=0.5$, $\beta_{W}=-0.2$ and $\beta_{Z}=0$.
  • Scenario B: MCAR data with 30% missingness;
    $W\sim$ Normal; $\alpha_{0}=0.1$ and $\alpha_{W}=\alpha_{X}=\beta_{W}=\beta_{Z}=0.5$.
  • Scenario C: MAR data with 30% missingness;
    $W\sim$ Normal; $\alpha_{0}=0.1$ and $\alpha_{W}=\alpha_{X}=\beta_{W}=\beta_{Z}=0.5$.
@@ 4.2 Simulation results
In Table 1, we present the results of the simulations for the three different scenarios (A-C).
@@
Method $\alpha_{0}$ $\alpha_{W}$ $\alpha_{X}$ $\alpha_{Z}$ $\beta_{W}$ $\beta_{X}$ $\beta_{Z}$
@@
Scenario D
Full data 0.03 (0.68) 0.93 0.02 (0.53) 0.95 0.06 (0.96) 0.95 0.06 (0.97) 0.96 0.01 (0.32) 0.96 0.03 (0.59) 0.94 0.03 (0.59) 0.95
Complete-case 0.37 (0.96) 0.41 0.05 (0.78) 0.93 0.16 (1.46) 0.95 0.14 (1.44) 0.95 0.01 (0.37) 0.94 0.03 (0.68) 0.94 0.04 (0.69) 0.95
Exact 0.04 (0.74) 0.94 0.02 (0.63) 0.98 0.06 (0.91) 0.94 0.06 (0.91) 0.94 0.01 (0.35) 0.92 0.03 (0.56) 0.93 0.03 (0.60) 0.94
Approximate 0.04 (0.79) 0.96 0.05 (0.86) 0.97 0.06 (1.00) 0.96 0.06 (1.01) 0.96 0.01 (0.38) 0.95 0.03 (0.59) 0.96 0.03 (0.59) 0.96
@@ 3 Empirical coverage of 95% confidence intervals
The performance of the imputation methods was assessed using different metrics: the mean squared error (MSE), the width of 95% confidence intervals (CI width), and the empirical coverage of 95% confidence intervals (Coverage). The full data scenario provides a baseline for comparison, as it describes these metrics for models fitted to the original data, without missing values.
@@
In the scenario with 15% MCAR (Scenario A), the complete-case analysis, obtained by discarding units with missing values for $W$ and then fitting a cure model on the complete data only, exhibits slightly higher MSE and wider confidence intervals compared to the full data scenario. As expected, estimates from the complete case analysis are not biased but have a larger variance since fewer observations are used. Both the exact and the approximate imputation methods performed similarly to the full data scenario, with only minor differences in CI widths, only for the estimates of the effects associated with $W$. In particular, we observe that multiple imputation lead to a decreased variance for the estimates of the parameters corresponding to fully observed variables.
@@
When the missingness is increased to 30%, still under a MCAR mechanism (Scenario B), this trend is confirmed: the complete-case analysis shows higher MSE and wider confidence intervals, while the imputation methods (via either the exact or approximate procedure) yield results close to the full data scenario. Although there are slight variations, these methods effectively preserve empirical coverage, with a slightly greater CI length when the effects of $W$ are considered.
@@
These results show that the proposed approaches provide reliable parameter estimates even in the presence of a moderate to high proportion of missing data. However, the true strength of our methods becomes evident in the MAR scenario (Scenario C). Here, the proposed imputation methods significantly outperform the complete-case analysis according to all evaluation metrics. This demonstrates that when data are missing at random, both exact and approximate imputation methods can effectively mitigate the impact of missing data, producing estimates close to those obtained with full data. In contrast, the complete-case analysis performs poorly, exhibiting a marked increase in MSE and a decrease in coverage as a result of the bias. Specifically, for $\alpha_{0}$, the MSE of the complete-case approach increased to 0.35 and the coverage decreased to 0.32. Additionally, the CI width for $\alpha_{X}$ was considerably wider than that obtained when using the imputation methods. Moreover, the estimate $\beta_{Z}$ shows a slight deterioration compared to scenarios A and B. Indeed, when using complete-case analysis, the missing data of $W$ leads to the removal of observations that might be critical for accurately estimating $\alpha_{X}$ and $\beta_{Z}$, especially if $W$ is correlated with $X$ and $Z$. This results in biased estimates and an increased MSE. In other words, by eliminating observations with missing $W$, the model’s ability to capture the true relationships involving $X$ and $Z$ may be reduced.
@@
The exact conditional distribution method outperforms the approximate method in terms of CI width, indicating more precise estimates. Across all scenarios, both the exact and the approximate methods consistently provide reliable estimates, maintaining the empirical coverage of the 95% confidence intervals. Particularly in cases where covariate $W$ is missing at random, the proposed methods significantly outperform the complete-case approach, demonstrating their ability to effectively handle complex missing data patterns.
4.3 Comparison with a full model
The simulation Scenario D corresponds to the method proposed by Beesley et al. (2016), which we will refer to as full model and serves as a benchmark for evaluating a further feature of the approach we propose here. In this scenario, we include in both components covariates whose impact on the outcome is null ($\alpha_{Z}=\beta_{X}=0$). In other words, we saturate the model by including irrelevant covariates in the incidence and the latency.
Thus, the set of parameters can be summarized as follows:
  • Scenario D: MAR data with 30% missingness;
    $W\sim$ Normal; $\alpha_{0}=0.1$, $\alpha_{W}=\alpha_{X}=\beta_{W}=\beta_{Z}=0.5$, and $\alpha_{Z}=\beta_{X}=0$
The comparison between Scenarios C and D highlights the advantages of a reduced model with selected covariates in each of the two parts over the full model discussed by Beesley et al. (2016). Focusing on the Exact and the Approximate methods, we observe consistent improvements in Scenario C across both the incidence and latency components. For instance, the MSE for $\alpha_{W}$ decreases from 0.02 in Scenario D to 0.01 in Scenario C using the exact method, and from 0.05 to 0.03 with the approximate method. This improvement is even more evident for $\alpha_{X}$, where the MSE decreases from 0.07 (Exact, Scenario D) to 0.04 (Exact, Scenario C), and from 0.07 to 0.05 for the approximate method.
In addition, confidence interval (CI) widths for all parameters are narrower in Scenario C. Specifically, the CI width for $\alpha_{W}$ improves from 0.65 (Exact, Scenario D) to 0.60 (Exact, Scenario C), and from 0.92 to 0.79 for the approximate method. A similar pattern is observed for $\beta_{W}$, where CI widths decrease from 0.36 (Exact, Scenario D) to 0.34 (Exact, Scenario C), and from 0.40 to 0.35 for the approximate method.
These results demonstrate the advantage of differentiating between covariates relevant to the incidence and to the latency component, as including irrelevant covariates leads to wider confidence intervals. In addition, we also observe that, if all covariates are used in both submodels when not necessary, the approximate method provides more reliable results compared to the exact approach. Even if the differences are modest in this simulation, since the number of parameters in the two models only differs by two, the benefits would be more substantial in scenarios with additional covariates, where the full model risks inefficiency due to overparameterization.
@@ 4.4 Algorithm performance under model misspecification
Appendix A and B provide the exact and approximate distributions for cases where the variable $W$ is included in only the incidence or the latency of the Cox PH cure model. However, we may lack information about the impact of $W$ on the two components. Therefore, we aim to study the performance of the proposed imputation algorithm when the model is misspecified, e.g. when $W$ is present in only one of the two components and we assume it is in both. To do this, we conduct a simulation study, examining two further scenarios:
  • Scenario E: MAR data with 30% missingness;
    $W\sim$ Normal; $\alpha_{0}=0.1$, $\alpha_{W}=\alpha_{X}=\beta_{Z}=0.5$, and $\beta_{W}=0$
  • Scenario F: MAR data with 30% missingness;
    $W\sim$ Normal; $\alpha_{0}=0.1$, $\alpha_{X}=\beta_{W}=\beta_{Z}=0.5$, and $\alpha_{W}=0$
These scenarios are similar to Scenario C introduced in section 4.1, but they differ in that the coefficient associated with the missing covariate $W$ is set to zero in either the latency (Scenario E, $\beta_{W}=0$) or the incidence component (Scenario F, $\alpha_{W}=0$). This allows us to test the algorithm’s ability to handle situations where the missing variable affects only one component of the Cox PH cure model. Note however, that these scenarios are not a comparison with Beesley et al. (2016) since the covariates $X$ and $Z$ are not the same.
Importantly, our focus is testing the algorithm’s performance under the MAR scenario. We expect that if the algorithm performs well in these more complex conditions, it will likely do so in scenarios where the missing data process is either less complex or it affects a lower portion of the data.
This simulation study provides insights into the robustness of the mixture cure model across the two distinct scenarios. In the previous setting, we discussed the inclusion of irrelevant variables without missing values in both model components, while here we investigate the effect of including a covariate with missing values also when it does not affect one of the submodels. Ideally, we would like a multiple imputation procedure that is robust to such model misspecification. Then, the model itself, through its fit to the available data, can suggest the appropriate placement of different covariates based on the significance of the coefficient estimates.
In Table 2, simulation results for scenarios E and F are shown. Analogously to what reported in Table 1, the performance of the imputation methods was assessed using the MSE, the CI width, and the empirical coverage of 95% confidence intervals.
@@
Method $\alpha_{0}$ $\alpha_{W}$ $\alpha_{X}$ $\beta_{W}$ $\beta_{Z}$
@@ 3 Empirical coverage of 95% confidence intervals
Quite similar to results in Table 1, the complete-case analysis shows a significant increase in MSE for $\alpha_{0}$, reaching 0.33 in Scenario E, with a corresponding coverage of only 0.31. This result reflects the method’s vulnerability to bias when missing data is not appropriately addressed. The complete-case method also displays increased CI widths, especially for $\alpha_{X}$, highlighting the loss of information and precision due to missing data.
The exact method generally maintains low MSE values in both scenarios and good coverage for most parameters. However, its performance sometimes falls short of the approximate method, as seen with $\alpha_{W}$ in Scenario E, where the exact method yields a coverage of 0.91 compared to 0.98 of the approximate method. The higher coverage of the approximate method can be attributed to its increased CI width, which provides a more conservative estimate, thereby encompassing more potential values of the parameter. The exact method’s performance again suggests it may be more sensitive to model assumptions, leading to increased bias when these assumptions are not met. Despite this, the results support conclusions about the general robustness of the model. The slightly poorer performance compared to the approximate method could be due to the tighter linkage between the missing data variable $W$ and the linear predictors, resulting in higher sensitivity to model assumptions and less flexibility in handling deviations.
The approximate method shows consistent performance with low MSEs in both scenarios and high coverage, highlighting its adaptability in dealing with model deviations. It achieves slightly better coverage for parameters like $\beta_{W}$ in Scenario F, reaching 0.94 compared to the exact method’s 0.90.
In Scenario F, the complete-case analysis exhibits a substantial increase in MSE values for $\alpha_{0}$ and poor coverage, confirming its inadequacy in scenarios with high levels of missingness. Conversely, the approximate method maintains reliable coverage and CI widths across parameters due to its adaptability and reduced reliance on strict assumptions.
Overall, these results suggest that the approximate method is particularly well-suited for real-world applications where knowledge of model structure is limited and data are affected by several sources of error. Its robust performance across scenarios E and F reinforces the idea that it is a viable alternative to the exact method.